OpenStudy (chycora):

6. (3 pt) Solve. p + 11 – 4p = 8 p =

OpenStudy (chycora):

i think it's 1.3

OpenStudy (anonymous):

is there options? lol k12

OpenStudy (anonymous):

It's 1 :) just substitute it for p in the equation!

OpenStudy (anonymous):

Its 1

OpenStudy (anonymous):

not suppose to give the answer @AMYCARTER

OpenStudy (chycora):

thanx guys

OpenStudy (welshfella):

p + 11 - 4p = 8 subtract 11 from both sides p +11 - 11- 4p = 8-11 -3p = -3 divide both sides by -3 p = 1